velue: 10.00 points Siegfried Basset is 65 years of age and has a life expectancy of 11 more years. He wishes to invest $20,700 in an annuity that will make a level payment at the end of each year until his death. If the interest rate is 8.0%, what income can Mr. Basset expect to receive each year? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Annual income Reterences Book & Resources 徘鲥割% 1d.sesegmes Worksheet Difmiculity: Intermediate Check my work
Answer #1

We have to get the Annual level payment which is received per year for 11 years and finally balance gets zero. So,

formula : Annual level Payment = P * i * (1+i)^n / (1+i)^n - 1
Annual level Payment = 20700 * 0.08 * (1+0.08)^11 / (1+0.08)^11 - 1
Annual level Payment = 1656 * (1.08)^11 / (1.08)^11 - 1
Annual level Payment = 1656 * 2.331639 / 2.331639 - 1
Annual level Payment = 3861.194184 / 1.331639
Annual level Payment = 2899.58 or 2900